Helium gas is contained in a tank with a pressure of 11.0 MPa. If the temperature inside the tank is 22.0 ºC and the volume of the tank is 15.8 L, determine the mass, in grams, of the helium in the tank.



1
Expert's answer
2022-04-07T04:06:17-0400

pV = nRT

p = 11.0 MPa = 11×106 Pa = 11×106 × 9.87×10-6 atm = 108.57 atm

T = 22.0°C = (22.0 + 273.15) K = 295.15 K

n = pV/RT = (108.57 atm)(15.8 L) / (0.0821 L atm/mol K)(295.15 K) = 70.79 mol

m(He) = n(He) × Mr(He) = 70.79 mol × 4.00 g/mol = 283.17 g